what is the period of the function y=sin(x/4)?
Answer by fractalier(6550)

The period of any function
y = a*sinBx is
2(pi) / B
so here
y = sin(x/4),
our B = 1/4, and the period is
2(pi) / (1/4) =
8(pi)